#1. Calculate the [OH−] of aqueous solution with the [H3O+] = 7.0×10−12 M .
Express your answer using two significant figures.

[OH−] = ???M

#2. Calculate the pH of aqueous solution with the pOH = 2.60.
Express your answer using two decimal places.

pH = ???
State whether the solution is acidic, basic, or neutral:

Acidic solution
Neutral solution
Basic solution
Thank you!!
1
Expert's answer
2015-11-19T12:25:16-0500
[ OH-] = 14.29x10^-4 M


Question:
#2. Calculate the pH of aqueous solution with the pOH = 2.60. Express your answer using two decimal places. State whether the solution is acidic, basic, or neutral

Solution:
pH + pOH = 14
If pOH = 2.60 than pH = 14 - pOH = 14 - 2.60 = 11.40

Answer:
pH = 11.40
The solution is basic